How to calculate theoritical NOx and SO2 (mg/Nm3) in gases at coal fired power plant, if known :

Coal Calorific Value = 4500 kcal/kg
Fuel consumption = 180 t/h
Excess air = 18%
Oxygen in exhaust gases = 5%
Flue Gas Temperature in exhaust gases = 135 deg C

Carbon Content = 30.110
Hydrogen Content = 3.610
Oxygen Content = 9.850
Sulfur Content = 0.330
Nitrogen Content = 0.700
Ash Content = 5.740
Moisture Content = 32.530

Thank you!
Answer #1


Theoritical No. and so (mg/Nm) in gases in wal fired power plant. given: tetonnes : fuel consumption 180 t/hr 180x10° leg Ih(Aus) The ne jure flue gas gas will contain CO, CO2, 4,0, 50,, Nz., NOXexcess Q (54.) reactions are stO₂- SO2 C+ to. - 2H+ +giveni s content = 0.330 mass of s in wal = 0.330 x 180x10² kg/hr : 59.4*10legiur. So, formed thmol)= 59.48103 kg Ihr 32.06how carbon content: Kalhr C: 30.110 x 180X10%, 5419.8 XIo Hydrogen content ! h = 3.610x 180 x 10² = 649.8 XI0² kg/hr content:The N₂ in in coming air that is unreacted the theoritical amt of 2 depends on il needed for combustion : complete clcmol) -oamt of flue gas in kmol + x 3.77 0₂ c 12.01 + h 2.016 + s 32.06 + n 28.02 + m 18.016 - = 5419.8 %10% + 649,8X10% + 1.05x10+5% of this is 5 x 558.87X10²: 27.94 X 10² kmol/hr excess o gas (including - the Amt of = 3591.14010² + 27.94X103 lcmol/hr Flo: Flowrate of flue gas (2) Q- 3619 kmol x 22.7 Nm3 : 82 153.19 Nms hr lemol lcmol . Now, amt of so = 1.85*10 ² lemol/hr. i 1Amt of NOxs assuming 20%. of N in wal.is converted to Nox. wal = 126 x 10² lcg. Amt of N₂ in to Nox this converts now 20% of